The segment joining (1, -3) and (4, -6 ) is extended a distance equal to one sixth of its own length. Find the terminal points.
Answer by greenestamps(13200)



I will assume that since the problem asks for the terminal points (plural), it means there are two answers -- one by extending the segment in either direction.

The difference in the x coordinates of the two points is 3; 1/6 of that is 1/2.

The difference in the y coordinates of the two points is also 3; 1/6 of that is again 1/2.

If starting at (1,-3) and going to (4,-6) and then going 1/6 of that distance more in that direction, then the coordinates of the terminal point are (4+1/2,-6-1/2) = (4.5,-6.5).

If starting at (4,-6) and going to (1,-3) and then going 1/6 of that distance more in that direction, then the coordinates of the terminal point are (1-1/2,-3+1/2) = (0.5,-2.5).

ANSWERS: (4.5, -6.5) or (0.5,-2.5)